1) A mass of 10 kg is inside a large hollow spherical mass (m = 2.0 x 1024 kg, r = 3.0 x 106 m). Its position is halfway between the center and the edge of the shell. Recall that G= 6.67 x 10-11 N m2 / kg2. Determine the gravitational force each exerts on the other.

A.148 N

B. 83 N

C. 74 N

D. 0 N

2) A mass of 10 kg is inside a large solid uniform spherical mass (m = 2.0 x 1024 kg, r = 3.0 x 106m). Its position is halfway between the center and the edge of the shell. Recall that G= 6.67 x 10-11 N m2/kg2. Determine the gravitational force each exerts on the other.

A.148 N

B. 83 N

C. 74 N

D. 0 N

3)

A mass of 10 kg is outside a large hollow spherical mass (m = 2.0 x 1024 kg, r = 3.0 x 106 m). Its position is 1.0 x 106 m above the surface of the shell. Recall that G= 6.67 x 10-11 N m2/kg2. Determine the gravitational force each exerts on the other.

A.148 N

B. 83 N

C. 74 N

D. 0 N
Answer #1

1. as we know shell exerts zero net force on object placed inside the shell

so using this fact we can say that force on object placed at halfway inside the shell is zero

so option D is correct

2. same question as question1

3. F = GMm/(r+h)^2

where M = 10kg h = 1 * 10^6 m

F = GMm/(r+h)^2 = 6.67* 10^(-11) *10 *2*10^24 */((3+1)*10^6)^2 = 83 N

OPTION B